OpenStudy (anonymous):

4.06] What is the equation of the line that passes through the point (2, 5) and has a slope of 2? y = 2x + 1 y = 2x + 3 y = 2x - 8 y = 2x - 3

OpenStudy (anonymous):

\(y=mx+b\) -> slope-intercept form pf a line where m is the slope, b is the y-intercept so you have m=2 and a point (2,5)=(x,y) y=2x+ b, solve for b using the points 5=2(2)+b b=1 so you have y=2x+1

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

welcome :)